# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: Present address: Microsoft Azure Quantum, Redmond, WA, USAthanks: Present address: Microsoft Azure Quantum, Redmond, WA, USA

### v1

@@ Abstract
To realize the potential of quantum computation, error-corrected machines are required [1] that can dramatically reduce the inevitable errors experienced by physical qubits. While significant progress has been made in approaching and exceeding the surface-code threshold in superconducting platforms [2, 3, 4], large gains in the logical error rate with increasing system size remain out of reach. This is due both to the large number of required physical qubits and the need to operate far below threshold. Importantly, by designing the error correction schemes to exploit the biases and structure of the physical errors, this threshold can be raised. Examples include codes that benefit from qubits with large Pauli-noise bias [5, 6] and erasure qubits [7, 8, 9, 10, 11], in which certain errors can be detected at the hardware level. Dual-rail qubits encoded in superconducting cavities are a promising erasure qubit wherein the dominant error, photon loss, can be detected and converted to an erasure. In these approaches, the complete set of operations, including two qubit gates, must be high performance and preserve as much of the desirable hierarchy or bias in the errors as possible. Here, we design and realize a novel two-qubit gate for dual-rail erasure qubits based on superconducting microwave cavities. The gate is high-speed ($\sim 500$ ns duration), and yields a residual gate infidelity after error detection below $0.1\%$, comparable to state-of-the-art results in any superconducting qubit platform. Moreover, we experimentally demonstrate that this gate largely preserves the favorable error structure of idling dual-rail qubits, making it an excellent candidate gate for enabling error correction and fault-tolerant computing. We measure low erasure rates of approximately $0.5\%$ per gate, as well as low and asymmetric dephasing errors that occur at least three times more frequently on control qubits compared to target qubits. Bit-flip errors are practically nonexistent and bounded at the few parts per million level. Such a high degree of error asymmetry has not been well explored but is extremely useful in quantum error correction and flag-qubit contexts, where it can create a faster path to error-corrected systems that rapidly suppress errors as they scale.
preprint: APS/123-QED
@@ I Introduction
A second way to improve performance is to first detect errors as they occur at the hardware level. For instance, if the dominant errors are leakage in the form of photon loss, as in photonic quantum computing [22, 23], the detectable loss of a qubit, called an erasure, can be very beneficial for error correction. For example, the well-known surface code has an increased threshold for phenomenological erasures of almost 25% per round [24, 25], and can additionally correct twice as many erasures as Pauli errors [26] for a given code distance. In neutral atoms, leakage of atoms to non-computational states can also be detected to convert them to erasures [11, 27], with similar benefits [10, 28] if other sources of errors remain small.
@@
Here, we propose and experimentally demonstrate a controlled-$Z$ (CZ) operation that preserves the error hierarchy in a system of superconducting cavity dual-rail qubits. This “SWAP-Wait-SWAP” (SWS) gate uses high-fidelity parametric operations to temporarily swap an excitation from one of the cavities (that comprise the dual-rail control qubit) to a transmon coupler connecting the two qubits, similar to the gate implementation in  [39]. By temporarily populating the coupler, we are able to utilize the strong (MHz level) dispersive shift between the coupler and one cavity of the other (target) dual-rail qubit. Upon swapping the excitation back into the control cavity, we realize an entangling CZ gate in $\sim 500$ nanoseconds.
Importantly, this excitation-preserving operation preserves the erasure-to-Pauli noise bias on both qubits, with erasure rates below 1%. After erasure detection, we find the residual dephasing errors are at the level of 0.1% per gate or below, as characterized by quantum state tomography and interleaved randomized benchmarking (RB). Moreover, we show that bit flips occur at less than a few parts per million per gate. Importantly, there is an additional error asymmetry wherein the target qubit erasures and dephasing rates are a factor of 3-4 times lower than those for the control qubit. This asymmetry can be beneficial when detecting error syndromes [40]. Finally, we study the error channel when one of the qubits suffers from photon loss errors during the gate, the dominant error in our system, and show both theoretically and experimentally that this amounts to a conditional-dephasing error on the unleaked qubit. All of these properties can be exploited by properly designed error correction codes where the performance levels we demonstrate are well past the predicted surface code thresholds [40, 11], and should thus allow significant logical error suppression with increasing surface code distance.
@@ II Results
Figure 1: Device hardware schematic and CZ gate sequence. a. Schematic of a two dual-rail cavity qubit system, similar to ref. [33], where the control and target qubit qubits are realized using pairs of three-dimensional $\lambda/4$ coaxial microwave cavities (red and blue). Three SQUID transmon couplers (grey and purple) each dispersively couple to pairs of cavity modes. These couplers are flux-pumped [35] to actuate parametric beamsplitter interactions for single qubit gates within the dual-rails, and also to implement the dual-rail two-qubit gate, which is achieved via the middle coupler. Each cavity is coupled to an additional transmon qubit (green) and readout resonator (blue lines) that are used for state preparation and measurement. b. Diagram of the SWS gate sequence between two dual-rail cavity qubits. This gate consists of three steps: (1) switching on the dispersive interaction by performing a SWAP operation between the control qubit ($a_{2}$ cavity) and the coupler $c$, (2) allowing the dispersive interaction between coupler and target qubit ($b_{1}$ cavity) to acquire phase and entangle the two dual-rail qubits, and (3) switch off the dispersive interaction by performing a SWAP between control qubit ($a_{2}$ cavity) and coupler, restoring population to the $a_{2}$ cavity.
@@
Our two-qubit gate makes use of a switchable dispersive interaction to implement a CZ gate between two dual-rail qubits composed of nominally linear microwave cavities. To explain our CZ gate, we first make a key observation: implementing an entangling interaction between the two nearest neighbor cavities results in entanglement between the two dual-rail qubits. Thus, only nearest-neighbor coupling is sufficient to implement the CZ gate. The physical system shown in Fig. 1a realizes this paradigm by utilizing a coupler $c$ that interacts dispersively with both the $a_{2}$ and $b_{1}$ cavities. For our CZ gate construction we utilize the dispersive interaction between the coupler and the target qubit, $\mathcal{H}_{\mathrm{int}}/\hbar=\chi_{bc}\hat{b}_{1}^{\dagger}\hat{b}_{1}\hat{c}^{\dagger}\hat{c}$, where $\chi_{bc}/2\pi=-1.51$ MHz in our system.
@@
Normally, this interaction is switched off as the coupler is an auxiliary mode, and nominally in its ground state. The interaction can be switched on by attempting to swap a photon between the cavity $a_{2}$ and coupler $c$, temporarily redefining the control dual-rail qubit between the non-local elements $a_{1}$ and $c$ (Fig. 1b). Such a swap here is implemented using a parametric sideband interaction between the coupler and cavity: $\mathcal{H}/\hbar=\frac{g_{ac}}{2}\left(\hat{a}_{2}^{\dagger}\hat{c}+\hat{a}_{2}\hat{c}^{\dagger}\right)$, similar to the interaction used to effect single-qubit gates in dual-rails. In our hardware, $g_{ac}/2\pi=4.23$ MHz. The dispersive interaction is then allowed to act for a time $t\sim\pi/\chi_{bc}$, thus entangling the two qubits. A final swap between the coupler and cavity $a_{2}$ returns the excitation back to the control cavity. The entangling phase $\phi_{e}=\chi_{bc}t_{\mathrm{wait}}$ has thus been imparted only to the |$\ket{01g10}$ initial state, resulting in a $\mathrm{CZ}$ gate.  [^1] Note that the phase of the two-qubit operation is parametrizable by adjusting the delay time, $t_{\mathrm{wait}}$. In this work, we always take $t\sim\pi/\chi_{bc}$. We have introduced the notation |$\ket{a_{1},a_{2},c,b_{1},b_{2}}$ to describe the full state of the system in the number basis, including both dual-rail qubits and the coupler mode.
@@
The construction of this two-qubit gate leads us to expect a simple error structure, which we observe and quantify below. This enables low error rates and gives rise to features that greatly benefit quantum error correction or flag-qubit error detection in quantum algorithms. First, the total number of excitations are preserved throughout the entire operation. That is to say, both control and target dual-rails never leave the single-excitation manifold (with a redefinition of the control dual rail during the gate operation to include the auxiliary coupler mode). This property ensures that any photon loss (even in the coupler) leads to a detectable error for our CZ gate. Second, this gate scheme also avoids detrimental effects due to cavity self-Kerr, as no mode is populated with more than one photon at any point during the gate. Third, errors on the target and control qubits can have significant asymmetry. This is due the fact that only the control qubit is temporarily stored (if cavity $a_{2}$ is initially excited) in the lossy coupling transmon. The control qubit thus experiences increased relaxation (with probability $p=t/T_{1}\sim 1\%$) and dephasing (with $p=t/T_{\phi}\sim 0.1\%)$, where we have assumed $T_{1}\sim 50~{}\mu$s and $T_{\phi}\sim 500~{}\mu$s for the typical coherence times of a transmon. In contrast, the target dual-rail is simply left in its idling state throughout, so it has a smaller probability of undergoing a detectable leakage ($p\simeq t/T_{\rm 1,cav}\sim 0.1\%$) and dephasing ($p\simeq t/T_{\rm\phi,DR}\sim 0.01\%$) given the idling single-photon decay times $T_{\rm 1,cav}\sim 500~{}\mu$s and dephasing times $T_{\rm\phi,DR}\sim 5$ ms. Fourth, we expect a strong suppression of bit-flip errors, as observed in the idling state of the dual-rails [33]. This is because the outer cavities $a_{1}$ and $b_{2}$ and their couplers remain untouched during the gate. Finally, we expect well-behaved propagation of leakage errors. The dominant leakage channel is decay to the vacuum state, which switches off the CZ-generating cross-Kerr interaction. Thus, if either of the dual-rails relax to vacuum, subsequent CZ gates simply do not occur (the operation is effectively the identity).
@@
This SWS gate can be calibrated via a straightforward procedure. In essence, it can be broken down into three pieces; i) swapping the photon from the control qubit cavity $a_{2}$ into the coupler $c$, ii) allowing the entangling phase from the dispersive interaction to accumulate over a specified delay time, and iii) swapping the photon back from the coupler into the control qubit cavity. In order to implement the CZ gate consisting of these pieces, we need to calibrate six parameters; the resonance frequency and duration of the parametric pump drive for cavity-coupler swap operation, the duration of the delay time needed to accumulate a $\pi$ entangling phase, the rotation angle for the second cavity-coupler swap to restore the photon population into the two-qubit manifolds, and the deterministic $Z$-phase of each single qubit gained during the gate. It is important to note that all these parameters can be independently calibrated in a sequential procedure; the one exception is the delay time, which requires prior adjustment of the rotation angle for the second cavity-coupler swap. A more detailed explanation of these parameters and the calibration protocol can be found in Appendix A.
@@
Figure 2: Benchmarking a dual-rail qubit CZ gate. a. Gate sequence used to quantify CZ gate fidelity via quantum state tomography. Applying a large number of odd number of repeated CZ gates ideally generates a Bell state, and increases our sensitivity to gate errors. For all cases except $N=1$, $X$ gates are applied to each qubit midway through the sequence, after the first $(N-1)/2$ CZ gates. This serves to echo out low-frequency dephasing noise and to eliminate effects of no-jump backaction. By changing the single qubit gates, $R_{c}$ and $R_{t}$, we can extract all 16 joint-Pauli correlators for the final state. b. Quantum state tomography after applying the gate sequence for $N=1$ CZ gates (dark) and $N=103$ CZ gates (light), represented by joint Pauli correlations. The ideal Bell state is defined by $\langle{II}\rangle=\langle{XX}\rangle=1$ and $\langle{YZ}\rangle=\langle{ZY}\rangle=-1$, with the residuals shown below. c. Repeated CZ gate Bell state tomography showing the increase in gate error with increasing gate repeats. We measure Bell state fidelity (upper panel, purple circles), state purity (upper panel, yellow squares), and postselected fraction, the fraction of shots for which no errors are detected (lower panel, grey), as a function of the number of CZ gates. Insets show data up to $N=15$ CZ gates. Error bars are generated from bootstrapped resampling of the data and are smaller than the size of the markers. d. Two-qubit interleaved RB showing reference (blue squares) and interleaved (purple circles) experiments. Each Clifford sequence is randomized 20 times and the mean and standard deviation over random seeds are plotted. The fraction of shots measured in the two-qubit codespace are plotted in the lower panel.
@@
Overall, our two-qubit system exhibits high fidelity SPAM and single-qubit gate fidelities advantageous for detailed characterization of our two-qubit gate. Single-qubit gates are performed by driving a beamsplitter interaction within the dual-rail for durations $\sim 200$ ns, achieving erasure rates of $\sim 0.1\%$ and post-selected RB error of $\sim 0.01\%$. By using end-of-the-line erasure-detected measurements, we are able to obtain low postselected SPAM of $\sim 0.02\%$ [33]. We provide additional results in Appendix C.
@@
Next, we perform repeated CZ gates to establish a more precise metric for the errors-per-CZ-gate, using the sequence shown in Fig. 2a. We apply an odd number ($N$) of total CZ gates to repeatedly entangle and disentangle the two qubits and subsequently characterize the resulting Bell state with quantum state tomography. We also include an $X$ gate after the first $(N-1)/2$ CZ gates which serves to eliminate the effect of no-jump backaction, first-order drift in control signal amplitudes, and any extraneous sources of low-frequency dephasing noise. We perform up to $N=103$ repeated gates and show the corresponding state tomography results in Fig. 2b, extracting an end-to-end postselected state fidelity of $87.4(1)\%$ and a state purity of $78.4(2)\%$. From this we infer, after post-selecting erasures and with no SPAM corrections, a bound of 0.12% on the remaining error-per-gate, among the best reported for superconducting qubits. The dependence on number of gates is shown in Fig. 2c, where we present state fidelity, state purity, and postselected fraction as a function of the number of repeated CZ gates. First, we can consider the behavior at lower gate repeats. As shown in the inset of Fig. 2c, our data shows a linear decrease in fidelity and purity out to $N=15$ repeated gates, from which we extract a post-selected infidelity of $0.03\%$ per gate. For larger gate depths, both the fidelity and purity of the final state display a quadratic decrease, which may be due to drift in calibration parameters, or fluctuations in the frequency of the coupling transmon on timescales of experimental shot averaging and will require further investigation. Adding more than one $X$ gate does not help to alleviate this effect. Finally, from our tomography data we also extract the postselected fraction, defined to be the fraction of experimental shots that were not assigned to an erasure. From fitting the results to an exponential decay, we extract a total erasure probability per CZ gate of $0.53(2)\%$.
@@
As an alternative method for determining the post-selected gate fidelity, we perform interleaved RB to test our CZ gate (Fig. 2d). RB in the presence of leakage errors sometimes requires taking into account the nuanced effects [42, 43] of leakage and seepage, where leaked states can silently re-enter the codespace. However, given our hardware error rates these effects are negligible. We use the gateset $\{X_{\frac{\pi}{2}},Z_{\pm\frac{\pi}{2}},Z,\mathrm{CZ}\}$ to generate the full two-qubit Clifford group, performing 20 randomized sequences for each Clifford sequence length. Because of the finite erasure probability ($0.69\%$) per gate, the usable sequence depth is limited to only a few hundred gates, as shown in the post-selected fraction curve in the lower panel. This complicates the fitting to the usual exponential form expected for the sequence yield (top panel of Fig. 3). We therefore use a linear fit to the curves for shorter depths (inset), obtaining an estimate for the error per CZ gate of $0.1\%$, in reasonably good agreement with the estimate from repeated Bell state tomography.
@@
Expected Measured
Erasure $\pi/\left(\chi_{bc}T_{1}^{\mathrm{c}}\right)\sim 0.5\%$ 10:1 10:1 $Z$-error on unleaked qubit
Dephasing $\pi/\left(\chi_{bc}T_{\phi}^{\mathrm{c}}\right)\sim 0.1\%$ 10:1 4:1 Commutes with gate; ancilla dephasing becomes stabilizer measurement error
@@
Table 1: Error hierarchy and asymmetry for the SWS gate. Values reflect both the expected and measured performance expected for a SWS gate and their effect on quantum error correction schemes.
@@
Having benchmarked the performance of the CZ gate, we now investigate the structure of gate errors on the control and target qubit. Because our CZ gate relies on swapping of the control qubit into the coupler and back, for part of the CZ gate, the control qubit will be physically defined between one cavity and the coupler. As such, gate errors, such as leakage and dephasing, will be largely set by the coupler coherences. On the other hand, the target qubit simply idles for the duration of the gate and so leakage and dephasing will be set by its longer idling coherences. With our system parameters we observe greater than 3 times lower errors on the target qubit than on the control qubit, which is a new and powerful property of our gate. For example, when performing error-correction one may choose the ancillae to always be the control qubits. Bit-flip errors are practically nonexistent in our CZ gate, which greatly limits the injection of errors into the data qubits of the logical qubit. In the following, we measure the error asymmetry for both leakage and dephasing errors as well as bit-flip errors.

### v2

# Bias-preserving and error-detectable entangling operations in a superconducting dual-rail system

@@ Abstract
To realize the potential of quantum computation, error-corrected machines are required [1] that can dramatically reduce the inevitable errors experienced by physical qubits. While significant progress has been made in reaching the surface-code threshold in superconducting platforms [2, 3, 4], large gains in the logical error rate with increasing system size still remain out of reach. This is due both to the large number of required physical qubits and the need to operate far below threshold. Importantly, by designing the error correction schemes to exploit the biases and structure of the physical errors, this threshold can be raised. Examples include codes that benefit from qubits with large Pauli-noise bias [5, 6] and erasure qubits [7, 8, 9, 10, 11], in which certain errors can be detected at the hardware level. Dual-rail qubits encoded in superconducting cavities are a promising erasure qubit wherein the dominant error, photon loss, can be detected and converted to an erasure. In these approaches, the complete set of operations, including two qubit gates, must be high performance and preserve as much of the desirable hierarchy or bias in the errors as possible. Here, we design and realize a novel two-qubit gate for dual-rail erasure qubits based on superconducting microwave cavities. The gate is high-speed ($\sim 500$ ns duration), and yields a residual gate infidelity after error detection below $0.1\%$, comparable to state-of-the-art results in any superconducting qubit platform. Moreover, we experimentally demonstrate that this gate largely preserves the favorable error structure of idling dual-rail qubits, making it an excellent candidate gate for enabling error correction and fault-tolerant computing. We measure low erasure rates of approximately $0.5\%$ per gate, as well as low and asymmetric dephasing errors that occur at least three times more frequently on control qubits compared to target qubits. Bit-flip errors are practically nonexistent and bounded at the few parts per million level. Such a high degree of error asymmetry has not been well explored but is extremely useful in quantum error correction and flag-qubit contexts, where it can create a faster path to error-corrected systems that rapidly suppress errors as they scale.
@@ I Introduction
A second way to improve performance is to first detect errors as they occur at the hardware level. For instance, if the dominant errors are leakage in the form of photon loss, as in photonic quantum computing [22, 23], the detectable loss of a qubit, called an erasure, can be very beneficial for error correction. For example, the well-known surface code has an increased threshold for phenomenological erasures of almost 25% per round [24, 25], and can additionally correct twice as many erasures as Pauli errors [26] for a given code distance. In neutral atoms and trapped ions, leakage to non-computational states can also be detected and converted to erasures [11, 27, 28], with similar benefits [10, 29] if other sources of errors remain small.
@@
Here, we propose and experimentally demonstrate a controlled-$Z$ (CZ) operation that preserves the error hierarchy in a system of superconducting dual-rail cavity qubits. This “SWAP-Wait-SWAP” (SWS) gate uses high-fidelity parametric operations to temporarily swap an excitation from one of the cavities (that comprise the dual-rail control qubit) to a transmon coupler connecting the two qubits, similar to the gate implementation in  [40]. By temporarily populating the coupler, we are able to utilize the strong (MHz level) dispersive shift between the coupler and one cavity of the other (target) dual-rail qubit. Upon swapping the excitation back into the control cavity, we realize an entangling CZ gate in $\sim 500$ nanoseconds.
Importantly, this excitation-preserving operation preserves the erasure-to-Pauli noise bias on both qubits, with erasure rates below 1%. After erasure detection, we find the residual dephasing errors are at the level of 0.1% per gate or below, as characterized by quantum state tomography and interleaved randomized benchmarking (RB). Moreover, we show that bit flips occur at less than a few parts per million per gate. Importantly, there is an additional error asymmetry wherein the target qubit erasures and dephasing rates are a factor of 3-4 times lower than those for the control qubit. This asymmetry can be beneficial when detecting error syndromes [41]. Finally, we study the error channel when one of the qubits suffers from photon loss during the gate, the dominant error in our system, and show both theoretically and experimentally that this amounts to a conditional-dephasing error on the unleaked qubit. All of these properties can be exploited by properly designed error correction codes where the performance levels we demonstrate are well past the predicted surface code thresholds [41, 11], and should thus allow significant logical error suppression with increasing surface code distance.
@@ II Results
Figure 1: Device hardware schematic and CZ gate sequence. a. Schematic of a two dual-rail cavity qubit system, similar to ref. [34], where the control and target qubits are realized using pairs of three-dimensional $\lambda/4$ coaxial microwave cavities (red and blue). Three SQUID transmon couplers (grey and purple) each dispersively couple to pairs of cavity modes. These couplers are flux-pumped [36] to actuate parametric beamsplitter interactions for single qubit gates within the dual-rails, and also to implement the dual-rail two-qubit gate, which is achieved via the middle coupler. Each cavity is coupled to an additional transmon qubit (green) and readout resonator (blue lines) that are used for state preparation and measurement. b. Diagram of the SWS gate sequence between two dual-rail cavity qubits. This gate consists of three steps: (1) switching on the dispersive interaction by performing a SWAP operation between the control qubit ($a_{2}$ cavity) and the coupler $c$, (2) allowing the dispersive interaction between coupler and target qubit ($b_{1}$ cavity) to acquire phase and entangle the two dual-rail qubits, and (3) switching off the dispersive interaction by performing a SWAP between control qubit ($a_{2}$ cavity) and coupler, restoring population to the $a_{2}$ cavity. c. Schematic illustration of photon population during the gate sequence for the four basis states.
@@
Our two-qubit gate makes use of a dispersive interaction to implement a CZ gate between two dual-rail qubits composed of nominally linear microwave cavities. To explain our CZ gate, we first make a key observation: implementing an entangling interaction between the two nearest neighbor cavities results in entanglement between the two dual-rail qubits. Thus, only nearest-neighbor coupling is sufficient to implement the CZ gate. The physical system shown in Fig. 1a realizes this paradigm by utilizing a coupler $c$ that interacts dispersively with both the $a_{2}$ and $b_{1}$ cavities. For our CZ gate construction we utilize the dispersive interaction between the coupler and the target qubit, $\mathcal{H}_{\mathrm{int}}/\hbar=\chi_{bc}\hat{b}_{1}^{\dagger}\hat{b}_{1}\hat{c}^{\dagger}\hat{c}$, where $\chi_{bc}/2\pi=-1.51$ MHz in our system.
@@
Normally, this static interaction term does not affect the system dynamics in any way, and is effectively ‘switched off’ since the coupler mode $c$ is in its ground state at the beginning of the gate. The interaction can be effectively ‘switched on’ by attempting to swap a photon between the cavity $a_{2}$ and coupler $c$, temporarily redefining the control dual-rail qubit between the non-local elements $a_{1}$ and $c$ (Fig. 1b). Such a swap here is implemented using a parametric sideband interaction between the coupler and cavity: $\mathcal{H}/\hbar=\frac{g_{ac}}{2}\left(\hat{a}_{2}^{\dagger}\hat{c}+\hat{a}_{2}\hat{c}^{\dagger}\right)$, similar to the interaction used to effect single-qubit gates in dual-rails. In our hardware, $g_{ac}/2\pi=4.23$ MHz. The dispersive interaction is then allowed to act for a time $t\sim\pi/\chi_{bc}$, thus entangling the two qubits. A final swap between the coupler and cavity $a_{2}$ returns the excitation back to the control cavity. The entangling phase $\phi_{e}=\chi_{bc}t_{\mathrm{wait}}$ has thus been imparted only to the |$\ket{01g10}$ initial state, resulting in a $\mathrm{CZ}$ gate  [^1] Note that the phase of the two-qubit operation is parametrizable by adjusting the delay time, $t_{\mathrm{wait}}$. In this work, we always operate near $t\sim\pi/\chi_{bc}$ to enact $\mathrm{CZ}$ gates.. We have introduced the notation |$\ket{a_{1},a_{2},c,b_{1},b_{2}}$ to describe the full state of the system in the number basis, including both dual-rail qubits and the coupler mode. Our gate construction allows us to effectively realize a large ‘switchable’ dispersive interaction between cavity modes $a_{2}$ and $b_{1}$, without needing to modulate the strength of the underlying static dispersive interactions in our system Hamiltonian. Instead, we are making use of a simple SWAP operation to an auxiliary mode in order to actuate our entangling interaction, which is much easier than directly engineering a tunable ‘$\chi$’ interaction in the system Hamiltonian [43].
@@
The construction of this two-qubit gate leads us to expect a simple error structure, which we observe and quantify below. This gate enables low error rates and gives rise to features that greatly benefit quantum error correction or flag-qubit error detection in quantum algorithms. First, the total number of excitations are preserved throughout the entire operation. That is to say, both control and target dual-rails never leave the single-excitation manifold (with a redefinition of the control dual rail during the gate operation to include the auxiliary coupler mode). This property ensures that any photon loss, even to the coupler mode, leads to a detectable error for our CZ gate. Second, this gate scheme also avoids detrimental effects due to cavity self-Kerr, as no mode is populated with more than one photon at any point during the gate. Third, errors on the target and control qubits can have significant asymmetry. This is due the fact that only the control qubit is temporarily stored (if cavity $a_{2}$ is initially excited) in the lossy coupling transmon. The control qubit thus experiences increased relaxation (with probability $p=t/T_{1}\sim 1\%$) and dephasing (with $p=t/T_{\phi}\sim 0.1\%)$, where we have assumed $T_{1}\sim 50~{}\mu$s and $T_{\phi}\sim 500~{}\mu$s for the typical coherence times of a transmon. In contrast, the target dual-rail is simply left in its idling state throughout, so it has a smaller probability of undergoing a detectable leakage ($p\simeq t/T_{\rm 1,cav}\sim 0.1\%$) and dephasing ($p\simeq t/T_{\rm\phi,DR}\sim 0.01\%$) given the idling single-photon decay times $T_{\rm 1,cav}\sim 500~{}\mu$s and dephasing times $T_{\rm\phi,DR}\sim 5$ ms. Fourth, we expect a strong suppression of bit-flip errors, as observed in the idling state of the dual-rails [34]. This is because the outer cavities $a_{1}$ and $b_{2}$ and their couplers remain untouched during the gate. Finally, we expect well-behaved propagation of leakage errors. The dominant leakage channel is decay to the vacuum state, which switches off the CZ-generating cross-Kerr interaction. Thus, if either of the dual-rails relax to vacuum, subsequent CZ gates simply do not occur (the operation is effectively the identity).
@@
This SWS gate can be calibrated via a straightforward procedure. In essence, it can be broken down into three pieces; i) swapping the photon from the control qubit cavity $a_{2}$ into the coupler $c$, ii) allowing the entangling phase from the dispersive interaction to accumulate over a specified delay time, and iii) swapping the photon back from the coupler into the control qubit cavity. In order to implement the CZ gate consisting of these pieces, we need to calibrate six parameters; the resonance frequency and duration of the parametric pump drive for cavity-coupler swap operation, the duration of the delay time needed to accumulate a $\pi$ entangling phase, the rotation angle for the second cavity-coupler swap to restore the photon population into the two-qubit manifolds, and the deterministic $Z$-phase of each single qubit gained during the gate. It is important to note that all these parameters can be independently calibrated in a sequential procedure; the one exception is the delay time, which requires prior adjustment of the rotation angle for the second cavity-coupler swap. A more detailed explanation of these parameters and the calibration protocol can be found in Appendix B.
@@
Figure 2: Benchmarking a dual-rail qubit CZ gate. a. Gate sequence used to quantify CZ gate fidelity via quantum state tomography. Applying an odd number $N$ of repeated CZ gates ideally generates a Bell state, and increases our sensitivity to gate errors. For all cases except $N=1$, $X$ gates are applied to each qubit midway through the sequence, after the first $(N-1)/2$ CZ gates. This serves to echo out low-frequency dephasing noise and to eliminate effects of no-jump backaction. By changing the single qubit gates, $R_{c}$ and $R_{t}$, we can extract all 16 joint-Pauli correlators for the final state. b. Quantum state tomography after applying the gate sequence for $N=1$ CZ gates (dark) and $N=103$ CZ gates (light), represented by joint Pauli correlations. The ideal Bell state is defined by $\langle{II}\rangle=\langle{XX}\rangle=1$ and $\langle{YZ}\rangle=\langle{ZY}\rangle=-1$, with the residuals shown below. c. Repeated CZ gate Bell state tomography showing the increase in gate error with increasing gate repeats. We measure Bell state fidelity (upper panel, purple circles), state purity (upper panel, yellow squares), and postselected fraction, the fraction of shots for which no errors are detected (lower panel, grey), as a function of the number of CZ gates. Insets show data up to $N=15$ CZ gates. Error bars are generated from bootstrapped resampling of the data and are smaller than the size of the markers. d. Two-qubit interleaved RB showing reference (blue squares) and interleaved (purple circles) experiments. Each Clifford sequence is randomized 20 times and the mean and standard deviation over random seeds are plotted. The fraction of shots measured in the two-qubit codespace are plotted in the lower panel.
@@
Overall, our two-qubit system exhibits high fidelity SPAM and single-qubit gate fidelities advantageous for detailed characterization of our two-qubit gate. Single-qubit gates are performed by driving a beamsplitter interaction within the dual-rail for durations $\sim 200$ ns, achieving erasure rates of $\sim 0.1\%$ and postselected RB error of $\sim 0.01\%$. By using end-of-the-line erasure-detected measurements, we are able to obtain low postselected SPAM errors of $\sim 0.02\%$ [34]. We provide additional results in Appendix A.
@@
Next, we perform repeated CZ gates to establish a more precise metric for the errors-per-CZ-gate, using the sequence shown in Fig. 2a. We apply an odd number ($N$) of total CZ gates to repeatedly entangle and disentangle the two qubits and subsequently characterize the resulting Bell state with quantum state tomography. We also include an $X$ gate after the first $(N-1)/2$ CZ gates which serves to eliminate the effect of no-jump backaction, first-order drift in control signal amplitudes, and any extraneous sources of low-frequency dephasing noise. We perform up to $N=103$ repeated gates and show the corresponding state tomography results in Fig. 2b, extracting an end-to-end postselected state fidelity of $87.4(1)\%$ and a state purity of $78.4(2)\%$. From this we infer, after postselecting erasures and with no SPAM corrections, a bound of 0.12% on the remaining error-per-gate, among the best reported for superconducting qubits. The dependence on number of gates is shown in Fig. 2c, where we present state fidelity, state purity, and postselected fraction as a function of the number of repeated CZ gates. First, we can consider the behavior at lower gate repeats. As shown in the inset of Fig. 2c, our data shows a linear decrease in fidelity and purity out to $N=15$ repeated gates, from which we extract a postselected infidelity of $0.03\%$ per gate. For larger gate depths, both the fidelity and purity of the final state display a quadratic decrease, which may be due to drift in calibration parameters, or fluctuations in the frequency of the coupling transmon on timescales of experimental shot averaging and will require further investigation. Adding more than one $X$ gate does not help to alleviate this effect. Finally, from our tomography data we also extract the postselected fraction, defined to be the fraction of experimental shots that were not assigned to an erasure. From fitting the results to an exponential decay, we extract a total erasure probability per CZ gate of $0.53(2)\%$.
@@
As an alternative method for determining the postselected gate fidelity, we perform interleaved RB to test our CZ gate (Fig. 2d). RB in the presence of leakage errors sometimes requires taking into account the nuanced effects [44, 45] of leakage and seepage, where leaked states can silently re-enter the codespace. However, given our hardware error rates these effects are negligible. We use the gateset $\{X_{\frac{\pi}{2}},Z_{\pm\frac{\pi}{2}},Z,\mathrm{CZ}\}$ to generate the full two-qubit Clifford group, performing 20 randomized sequences for each Clifford sequence length. Because of the finite erasure probability ($0.69\%$) per gate, the usable sequence depth is limited to only a few hundred gates, as shown in the postselected fraction curve in the lower panel. This complicates the fitting to the usual exponential form expected for the sequence yield (top panel of Fig. 2d). We therefore use a linear fit to the curves for shorter depths (inset), obtaining an estimate for the error per CZ gate of $0.1\%$, in reasonably good agreement with the estimate from repeated Bell state tomography.
@@
Typical Measured
Erasure $\pi/\left(\chi_{bc}T_{1}^{\mathrm{c}}\right)\sim 0.5\%$ 10:1 4:1 $Z$-error on unleaked qubit
Dephasing $\pi/\left(\chi_{bc}T_{\phi}^{\mathrm{c}}\right)\sim 0.1\%$ 10:1 3.5:1 Commutes with gate; ancilla dephasing becomes stabilizer measurement error
@@
Table 1: Error hierarchy and asymmetry for the SWS gate. Values reflect both the typical and measured performance expected for a SWS gate and their effect on quantum error correction schemes.
@@
Having benchmarked the performance of the CZ gate, we now investigate the structure of gate errors on the control and target qubit. Because our CZ gate relies on swapping of the control qubit into the coupler and back, for part of the CZ gate, the control qubit will be physically defined between one cavity and the coupler. As such, gate errors, such as leakage and dephasing, will be largely set by the coupler coherences. On the other hand, the target qubit simply idles for the duration of the gate and so leakage and dephasing will be set by its longer idling coherences. With our system parameters we observe about 3.5 times lower errors on the target qubit than on the control qubit. Counterintuitively, this error asymmetry can be a highly beneficial property, particularly in an error-correction context. For example, when performing error-correction one may assign the ancilla (data) qubit to always correspond to the control (target) qubit. With this choice, data qubits will suffer lower dephasing errors, and ancilla qubits errors will just lead to syndrome measurement errors. Furthermore, bit-flip errors are practically nonexistent in our CZ gate, which greatly limits the injection of errors into the data qubits of the logical qubit. In the following, we measure the error asymmetry for both leakage and dephasing errors as well as bit-flip errors.